# Exploring the Naturalness of AI-Generated Images

### 14 Dec 2023 v2

@@ 4 The Proposed JOINT and JOINT++
Studies in neurosciences [33, 22, 67] suggest that humans possess two distinct visual systems, which follow two main pathways, i.e., the dorsal stream and ventral stream, to handle low-level and high-level visual perception, respectively. To align model behavior with human perception process, we propose the Joint Objective Image Naturalness evaluaTor (JOINT) that models human naturalness reasoning by simultaneously considering the impact of technical and rationality perspectives, shown in Fig. 7.
Specifically, the technical perspective mainly refers to the low-level distortions, e.g., texture losses, blurs, and artifacts [89, 88, 81, 1], while the rationality of image is largely related to high-level visual information including semantic, attribute theme, and layout that are also the key elements of aesthetic assessment [29, 28, 49, 95]. Notably, some of the perceptual factors are related to both perspectives, such as contrast that belongs to technical distortion and can affect color harmonization, or artifacts (a generative technical distortion but highly coupled with the existence and layout of the image contents). To allocate two branches with corresponding learning interests, we present several specific designs (e.g., patch partition, feature regularization) and two different training schemes, illustrated as follows.
@@ 4.1 The Technical Prior Branch
For technical prior branch, we explicitly guide the model to prioritize the technical distortions while minimizing the impact of semantic information. Concretely, we randomly crop the image $\mathcal{I}$ into size-fixed patches and stitch them together ($\mathcal{I}_{rand}$) to disorganize most contents and layout while retaining technical distortions, thus destroying semantic information and rationality factors in images [87, 88]. However, different from most global technical distortion, generated artifacts could become unrecognizable by random patch partition. Therefore, we localize possible perceptual artifacts first and bypass these regions to keep their local distortion information.
Perceptual Artifacts-guided Patch Partition. Numerous research efforts have sought to localize the edited regions [110, 86], which basically involves training a model to pinpoint systematic inconsistencies in generated images. More recently, Zhang et al. [102, 109] further expand this task to a fine-grained level that not only predicts inpainted areas but also detects and segments artifact areas that are noticeable to human perception. In this work, we use the detection model provided in [109] as the artifacts extractor to guide the patch partition. Note that we do not focus on the detection process in this work. Specifically, given an image $\mathcal{I}$ of size $\mathcal{I}_{W}\times\mathcal{I}_{H}$, the patch partition can be formulated as:
@@ 4.2 The Rationality Perceiving Branch
Since the high-level semantic information of rationality concerns is likewise of interest to the image aesthetic assessment (IAA), we pre-train this branch first with IAA database and introduce a deep feature regularization to mitigate the effect of technical quality.
@@
4.3 Learning Objectives
We have discussed in Sec. 3.3 that the overall naturalness score $\mathrm{MOS}$ can be approximated as a weighted sum of $\mathrm{MOS_{T}}$ and $\mathrm{MOS_{R}}$. Here, we propose an indirect supervision ($\mathcal{L}_{\mathrm{IS}}$) scheme to optimize two branches using the overall naturalness $\mathrm{MOS}$. However, the subjective bias in two perspectives can cause large absolute prediction errors and reduce the prediction accuracy for each branch. Inspired by [48, 46, 88], we add the Spearman Rank-order Correlation Coefficient (SRCC) loss as a restraint to boost the prediction monotonicity of models. Overall, JOINT learns to assess image naturalness by jointly optimizing:
${\mathcal{L}_{{\mathrm{C}}}}={\mathcal{L}_{{\mathrm{MSE}}}}+\alpha{\mathcal{L}_{{\mathrm{SRCC}}}},\$ (4)
Subjective Weighting Strategy. According to the subjective study in AGIN, we adopt a simple but effective weighting strategy to compute the overall naturalness prediction ($\hat{S}_{\mathrm{N}}$) from two perspectives: $\hat{S}_{\mathrm{N}}=0.145\hat{S}_{\mathrm{T}}+0.769\hat{S}_{\mathrm{R}}$. With better performance in experiments, this strategy further demonstrates the observations in Sec. 3.3.
5 Experiments
In this section, we conduct experiments to verify: (1) the necessity of AGIN database (Sec. 5.2). (2) can JOINT and JOINT++ better assess the overall naturalness of AGIs (Sec. 5.3)? Moreover, ablation studies are included (Sec. 5.4).
5.1 Experimental Settings
@@ 5.3 Evaluation on the AGIN
Figure 8: Qualitative Studies of JOINT and JOINT++. Visualizations of images in the AGIN that have opposite technical quality and rationality. Zoom-in for better visualization.
@@
TANet [29] 0.5367 0.5587 0.4731 0.4762 0.4782 0.4535
Del. Transf. [28] 0.5882 0.6134 0.5037 0.4942 0.4805 0.4961
SAAN [95] 0.4299 0.4380 0.4009 0.4160 0.4196 0.4184
@@
w/o Localization 0.811 0.816 0.755 0.769 0.782 0.794
w/o Regularization 0.814 0.820 0.729 0.738 0.758 0.766
w/o Multi-perspective 0.768 0.781 0.703 0.712 0.727 0.733
@@
Table 4: Ablation study of JOINT (I): the effects of specific designs in technical and rationality branches.
@@
$\checkmark$ 0.817 0.824 0.720 0.724 0.725 0.744
$\checkmark$ 0.687 0.699 0.756 0.771 0.767 0.763
$\checkmark$ $\checkmark$ 0.753 0.768 0.732 0.744 0.759 0.755
$\checkmark$ $\checkmark$ $\checkmark$ 0.711 0.723 0.746 0.762 0.799 0.803
@@
Table 6: Ablation study of JOINT++: the learning objectives.
5.4 Ablation Studies
Effects of Specific Designs. In Tab. 4, we verify the effect of three special designs in JOINT by keeping other parts the same. First, it is superior to the variant w/o Localization that randomly shuffles the patches and destruct the perceptual artifact regions, showing the necessity of preserving the local artifact distortion information. Secondly, with the deep feature regularization, JOINT is able to focus more on the rationality perspective. Furthermore, JOINT is notably better than the variant w/o Multi-perspective that directly takes the original images as inputs of both branches, proving the effectiveness of multi-perspective joint learning strategy.
Effects of Subjective Weighting Strategy. We discuss the subjective weighting strategy in Tab. 5. It can be observed that any single branch can not adequately represent the naturalness, and directly taking $\hat{S}_{\mathrm{T}}+\hat{S}_{\mathrm{R}}$ as overall naturalness without weighting will also bring a notable performance decrease ($-5.01\%$/$-5.98\%$ in terms of SRCC/PLCC), which further supports the observations found in the AGIN.
Effects of Learning Objectives. As shown in Tab. 6, supervised by corresponding MOS labels yields more accurate predictions. Compared with $\mathcal{L}_{\mathrm{IS}}$, $\mathcal{L}_{\mathrm{FS}}$ achieves around $+1.71\%$/$+1.21\%$, $+3.97\%$/$+2.98\%$, and $+2.50\%$/$+3.11\%$ performance gains in terms of technical, rationality, and naturalness assessment, respectively. It is also worth noting that combining $\mathcal{L}_{\mathrm{IS}}$ with $\mathcal{L}_{\mathrm{FS}}$ can significantly improve the prediction accuracy of overall naturalness. Additionally, even without MOS labels for each perspective, $\mathcal{L}_{\mathrm{IS}}$ can still achieve comparable performance, which suggests the feasibility of modeling the mechanisms of human perception of naturalness using both perspectives.
@@ 6 Conclusion
In this paper, we contribute AGIN database and the first subjective evaluation aimed at exploring the impact of technical and rationality perspectives on the naturalness of AGIs. Besides, we propose JOINT and JOINT++, the objective naturalness evaluators that achieve higher alignment with human opinions against existing IQA and IAA approaches. Our work benefits community by 1) presenting AGIN, which enables research on benchmarking and evaluating naturalness of AGIs by multi-dimensional human ratings; 2) encouraging new research on the naturalness assessment of AGIs via analysis of technical and rationality features; 3) promoting the development of better INA algorithms for AGIs or other forms of AI-generated multimedia.
Database Image Source #Content #Image Perspective Distortion
LIVE (2004) [78] Kodak test set 30 779 Quality 5 Artificial
TID2008 (2008)[72] Kodak test set 25 1,700 Quality 17 Artificial
CSIQ (2009)[43] Kodak test set 30 866 Quality 6 Authentic
TID2013 (2013)[71] Kodak test set 25 3,000 Quality 24 Artificial
LIVEC (2015)[21] Camera 1162 1162 Quality 15 Authentic
WED (2017)[61] Internet 4,744 94,880 Quality 4 Artificial
MDID (2017)[82] Internet 20 1,600 Quality 5 Artificial
KADID-10k (2019)[53] Internet 81 10,125 Quality 25 Artificial
KonIQ-10k (2020)[32] Multimedia 10,073 10,073 Quality -- Authentic
AGIN (Ours) AI-generated 6,049 6,049 Technical+Rationality+Naturalness 18 Generative
Figure 9: Additional results on generated images in AGIN database. We show four examples for each model in five tasks (text-to-image, image translation, image inpainting, image colorization, and image editing). Zoom-in for better viewing.
@@
[…]
Appendix
Appendix A Extended Details of AGIN Database
In this section, we provide additional details for the construction of the AGIN database. We define the task of evaluating the naturalness of AI-generated images as a special case of image quality assessment (IQA). As shown in Tab. 7, AGIN differs from the existing IQA database in three aspects: image source, evaluation perspective, and distortion type. Previous IQA databases mainly focus on post-capture artificial distortion (e.g., masked noise, JPEG compression) or in-capture authentic distortions (e.g., motion blur, exposure), while AI-generated images possess entirely different generative patterns and richer contents, posing new challenges for constructing a comprehensive database. Since we focus on the naturalness assessment of AI-generated images, only prompts that describe objective objects or phenomena and source images with authentic style are used to build the database. It is unnecessary to add abstract or surreal concepts, which are prone to appear unnatural. To cover diverse appearances of naturalness issues, we collect multi-sourced images from five tasks, as depicted in Tab. 8. Exemplar images for each model are shown in Fig. 9.
A.1 Collecting AI-generated Images
@@
Detailed Information of Image Editing Models. There exists a huge demand for providing flexible and controllable image editing means in daily life, ranging from individual users to professional applications. Improper instruction may cause unnaturalness between the edited areas and surrounding contents. Here, we choose the well-known text-guided image editing model (InstructPix2Pix [6]) and an interactive model (DragGAN [69]) to generate natural or unnatural images. For InstructPix2Pix, we apply text instructions, such as “add sth to sth”, “turn it into sth”, “replace sth the with sth”, to edit images. The source images are from ADE20K [113] and Landscape [13] datasets. For DragGAN, we choose images from the DeepFashion [57] dataset and set up two anchor points with certain drag directions. Then, DragGAN will move the handle point to reach its corresponding target point, thus completing the image editing procedure. Moreover, we extract 519 images from MagicBrush [100], a large-scale, manually annotated dataset for instruction-guided real image editing. Since it already covers diverse editing scenarios, we do not re-process these images.
Figure 14: Score distributions for male and female with different categories. From left to right are the violin plots of technical quality, rationality, and overall naturalness scores.
A.2 Detailed Information of Human Evaluation
@@
A.3 Detailed Score Distribution
Detailed score distribution of different categories. As shown in Fig. 13, we visualize the detailed score distribution of different categories for all participants. It can be observed that the real images in AGIN have relatively high technical, rationality, and naturalness scores, which is consistent with our intention of controlling the quality of the rating results. Besides, AGIs from the image translation task have the widest distribution of scores for technical, rationality, and naturalness. In general, regardless of the number of images, all categories have a relatively balanced score distribution in terms of technical quality, rationality, and naturalness.
Detailed score distribution of different categories in terms of gender. We also visualize the detailed score distribution of different categories for males and females in Fig. 14. We can observe that the score distribution of different categories for males and females is basically consistent except for the category of text-to-image. We speculate that people may have cognitive differences for familiar or unfamiliar objects.
@@ A.3 Detailed Score Distribution
Figure 16: Comparison of the original image and its corresponding low-frequency map from the AGIN database. Zoom-in for better visualization.
Appendix B Detailed Structure of the JOINT
@@ B.1 Low-Frequency Map
B.2 Definition of Wasserstein Distance
@@ B.3 Training Objective
Appendix C More Experimental Details
C.1 Implementation Details
@@ C.2 Evaluation Metrics
Figure 17: Success case (I) of JOINT. The images possess better rationality but worse technical quality (e.g., blur, detail, contrast).
Figure 18: Success case (II) of JOINT. The images possess better rationality but worse technical quality (e.g., artifacts, blur, detail).
Figure 19: Success case (III) of JOINT. The images possess better technical quality but worse rationality (e.g., existence, layout, context).
Figure 20: Failure case of JOINT. The images have better technical quality but worse rationality (e.g., existence, context, layout).
C.3 Extended Qualitative Results
Success Cases. We visualized several successful cases (i.e., when the predicted results are consistent with subjective annotators in each dimension) for the proposed JOINT in Fig. 17, Fig. 18, and Fig. 19. Specifically, we can observe that our JOINT is sensitive to global technical distortions such as blur, lack of detail, and contrast (Fig. 17). Moreover, the perceptual artifacts-guided patch partition strategy endows JOINT with the ability to measure the severity of local artifacts (Fig. 18). On the contrary, the rationality perceiving branch is not sensitive to technical distortion since we add deep feature regularization with filtered low-frequency information, and it is very sensitive to the rationality of contents and can recognize the very unusual compositions and objects in images (Fig. 19). All these cases further demonstrate the effectiveness of the JOINT in jointly learning from both perspectives in image naturalness assessment.
Failure Cases. As for the failure cases, we notice that most of them are difficult to recognize (poor sensory clarity, the top one in Fig. 20) or are extremely blurry with a small part of clear areas (mirror reflections, the bottom one in Fig. 20). Such cases are more frequent in photography that photographers gain prominence over the target by blurring the background, yet the model can experience challenges in predicting technical quality and rationality, resulting in a failure in naturalness evaluation.

### 04 Mar 2024 v3

@@ 4 The Proposed JOINT and JOINT++
Studies in neurosciences [31, 20, 66] suggest that humans possess two distinct visual systems, which follow two main pathways, i.e., the dorsal stream and ventral stream, to handle low-level and high-level visual perception, respectively. To align model behavior with human perception process, we propose the Joint Objective Image Naturalness evaluaTor (JOINT) that models human naturalness reasoning by simultaneously considering the impact of technical and rationality perspectives with two independent branches, shown in Fig. 7. To allocate these two branches with corresponding learning interests, we present several specific designs (e.g., patch partition, feature regularization) and two different training schemes, illustrated as follows.
@@ 4.1 The Technical Prior Branch
For technical prior branch, we explicitly guide the model to prioritize the technical distortions while minimizing the impact of semantic information. Concretely, we randomly crop the image $\mathcal{I}$ into size-fixed patches and stitch them together ($\mathcal{I}_{rand}$) to disorganize most contents and layout while retaining technical distortions, thus destroying semantic information and rationality factors in images [89, 90]. However, different from most global technical distortion, generated artifacts could become unrecognizable by random patch partition. Therefore, we propose to localize possible perceptual artifacts first and bypass these regions to keep their local distortion information.
Perceptual Artifacts-guided Patch Partition. Numerous research efforts have sought to localize the edited regions [111, 88], which basically involves training a model to pinpoint systematic inconsistencies in generated images. More recently, Zhang et al. [102, 110] further expand this task to a fine-grained level that not only predicts inpainted areas but also detects and segments artifact areas that are noticeable to human perception. In this work, we use the detection model proposed in [110] as the artifacts extractor to guide the patch partition. Note that we do not focus on the detection process in this work. Given an image $\mathcal{I}$ of size $\mathcal{I}_{W}\times\mathcal{I}_{H}$, the patch partition can be formulated as:
@@ 4.2 The Rationality Perceiving Branch
Since the high-level semantic information in rationality concerns is likewise of interest to the image aesthetic assessment (IAA), we pre-train this branch first with an IAA database and introduce a deep feature regularization to mitigate the effect of technical quality.
@@
4.3 Learning Objectives
We propose to optimize two branches using the overall naturalness $\mathrm{MOS}$ via indirect supervision ($\mathcal{L}_{\mathrm{IS}}$). However, the subjective bias between two perspectives can cause large absolute prediction errors and reduce the prediction accuracy for each branch. Hence, we add the Spearman Rank-order Correlation Coefficient (SRCC) loss [47, 45, 90] as a restraint to boost the prediction monotonicity of models. Overall, JOINT learns to assess image naturalness by minimizing:
${\mathcal{L}_{{\mathrm{C}}}}={\mathcal{L}_{{\mathrm{MSE}}}}$ $+\alpha{\mathcal{L}_{{\mathrm{SRCC}}}},$ (5)
Subjective Weighting Strategy. According to the subjective study in AGIN, we adopt a simple but effective weighting strategy to compute the overall naturalness prediction ($\hat{S}_{\mathrm{N}}$) from two perspectives: $\hat{S}_{\mathrm{N}}=0.145\hat{S}_{\mathrm{T}}+0.769\hat{S}_{\mathrm{R}}$. With better performance in experiments (Tab. 6), this strategy further demonstrates the observations in Sec. 3.3.
5 Experiments
5.1 Experimental Settings
@@ 5.3 Evaluation on the AGIN
[…]
BRISQUE [62]
0.4867
0.4909
0.3608
0.3684
0.3745
0.4067
NIQE [63]
0.4235
0.4279
0.3144
0.3211
0.3358
0.3378
[…]
0.7623
0.7661
0.6834
0.6838
0.7057
0.7128
HyperIQA [82]
0.7752
0.7806
0.7196 0.7292
0.7365
0.7509
MUSIQ [37]
0.7286
0.7355
0.6974
0.7013
0.7066
0.7103
UNIQUE [108]
0.7358
0.7434
0.6583
0.6685
0.6772
0.6789
MANIQA [95] 0.7763 0.7817
0.7192
0.7217
0.7385
0.7343
[…]
0.4763
0.4833
0.4532
0.4596
0.4483
0.4528
TANet [27]
0.5367
0.5587
0.4731
0.4762
0.4782
0.4535
Del. Transf. [26]
0.5882
0.6134
0.5037
0.4942
0.4805
0.4961
SAAN [96]
0.4299
0.4380
0.4009
0.4160
0.4196
0.4184
@@
Figure 8: Qualitative Studies of JOINT and JOINT++. Visualizations of images that have opposite technical quality and rationality. Zoom-in for better visualization.
@@
Table 5: Ablation study of JOINT (I): the effects of specific designs.
@@
[…]
w/o Localization
0.811
0.816
0.755
0.769
0.782
0.794
w/o Regularization
0.814
0.820
0.729
0.738
0.758
0.766
w/o Multi-perspective
0.768
0.781
0.703
0.712
0.727
0.733
@@
[…]
$\checkmark$
0.817 0.824
0.720
0.724
0.725
0.744
$\checkmark$
0.687
0.699
0.756 0.771
0.767
0.763
$\checkmark$
$\checkmark$
0.753
0.768
0.732
0.744
0.759
0.755
$\checkmark$
$\checkmark$
$\checkmark$
0.711
0.723
0.746
0.762
0.799 0.803
5.4 Ablation Studies
Effects of Specific Designs. In Tab. 5, we verify the effect of three special designs in JOINT by keeping other parts the same. First, it is superior to the variant w/o Localization that randomly shuffles the patches and destructs the perceptual artifact regions, showing the necessity of preserving the local artifact distortion information. Secondly, with the deep feature regularization, JOINT is able to focus more on the rationality perspective. Furthermore, JOINT is notably better than the variant w/o Multi-perspective that directly takes the original images as inputs of both branches, proving the effectiveness of multi-perspective joint learning strategy.
Effects of Subjective Weighting Strategy. As reported in Tab. 6, any single branch can not adequately represent the naturalness, and directly taking $\hat{S}_{\mathrm{T}}+\hat{S}_{\mathrm{R}}$ as overall naturalness without weighting will also bring a notable performance decrease ($-5.01\%$/$-5.98\%$ in terms of SRCC/PLCC), which further supports the observations found in the AGIN.
Effects of Learning Objectives. As shown in Tab. 7, supervised by corresponding MOS labels yields more accurate predictions. Compared with $\mathcal{L}_{\mathrm{IS}}$, $\mathcal{L}_{\mathrm{FS}}$ achieves around $+1.71\%$/$+1.21\%$, $+3.97\%$/$+2.98\%$, and $+2.50\%$/$+3.11\%$ performance gains in terms of technical, rationality, and naturalness assessment, respectively. It is also worth noting that combining $\mathcal{L}_{\mathrm{IS}}$ with $\mathcal{L}_{\mathrm{FS}}$ can significantly improve the prediction accuracy of overall naturalness. Additionally, even without MOS labels for each perspective, $\mathcal{L}_{\mathrm{IS}}$ can still achieve comparable performance, which suggests the feasibility of modeling the human perception of naturalness from technical and rationality perspectives.
Table 7: Ablation study of JOINT++: the learning objectives.
@@ 6 Conclusion
In this paper, we contribute the AGIN database and the first subjective evaluation aimed at exploring the impact of technical and rationality perspectives on the naturalness of AGIs. Besides, we propose JOINT, an objective naturalness evaluator that achieves higher alignment with human opinions against existing IQA and IAA approaches. Our work benefits the community by 1) presenting AGIN, which enables research on benchmarking and evaluating the naturalness of AGIs by multi-dimensional human ratings; 2) encouraging new research on the naturalness assessment of AGIs via analysis of technical and rationality features; 3) promoting the development of better INA algorithms for AGIs or other forms of AI-generated multimedia.
In this supplementary material, we provide additional details that we do not mention in the main paper.
Appendix 0.A Extended Details of AGIN Database
In this section, we provide additional details for the AGIN database. We define the task of evaluating the naturalness of AI-generated images as a special case of image quality assessment (IQA). As shown in Tab. 1 from the main paper, AGIN differs from the existing IQA database in three aspects: image source, evaluation perspective, and distortion type. Previous IQA databases mainly focus on post-capture artificial distortion (e.g., masked noise, JPEG compression) or in-capture authentic distortions (e.g., motion blur, exposure), while AI-generated images possess entirely different generative patterns and richer contents, posing new challenges for constructing a comprehensive database. The proposed AGIN provides guidance for redefining image quality in the AIGC field. Note that since we focus on the naturalness assessment of AI-generated images, only prompts that describe objective objects or phenomena and source images with authentic style are used to build the database. It is unnecessary to add abstract or surreal concepts, which are prone to appear unnatural. To cover diverse appearances of naturalness issues, we collect multi-sourced images from five tasks, as depicted in Tab. 8. Exemplar images for each model are shown in Fig. 9.
Figure 9: Additional results on generated images in AGIN database. We show four examples for each model in five tasks (text-to-image, image translation, image inpainting, image colorization, and image editing).
[…]
@@ Appendix 0.A Extended Details of AGIN Database
0.A.1 Collecting AI-generated Images
@@
Detailed Information of Image Editing Models. There exists a huge demand for providing flexible and controllable image editing means in daily life, ranging from individual users to professional applications. Improper instruction may cause unnaturalness between the edited areas and surrounding contents. Here, we choose the well-known text-guided image editing model (InstructPix2Pix [6]) and an interactive model (DragGAN [68]) to generate natural or unnatural images. For InstructPix2Pix, we apply text instructions, such as “add sth to sth", “turn it into sth", “replace sth the with sth", to edit images. The source images are from ADE20K [114] and Landscape [11] datasets. For DragGAN, we choose images from the DeepFashion [56] dataset and set up two anchor points with certain drag directions. Then, DragGAN will move the handle point to reach its corresponding target point, thus completing the image editing procedure. Moreover, we extract 519 images from MagicBrush [101], a large-scale, manually annotated dataset for instruction-guided real image editing. Since it already covers diverse editing scenarios, we do not re-process these images.
@@
0.A.2 Detailed Information of Human Evaluation
@@
Figure 14: More comparisons of image generation models. We visualize the averaged MOS for technical, rationality, and overall naturalness perspectives.
Figure 15: Score distributions for male and female subjects with different categories.
0.A.3 Detailed Score Distribution
Detailed score distribution of different categories. As shown in Fig. 13, we visualize the detailed score distribution of different categories for all participants. It can be observed that the real images in AGIN have relatively high technical, rationality, and naturalness scores, which is consistent with our intention of controlling the quality of the rating results. Besides, AGIs from the image translation task have the widest distribution and lowest average scores for technical, rationality, and naturalness. We can roughly observe from Fig. 14 that the naturalness and its two perspectives of AGIs have improved over the years. In general, regardless of the number of images, all categories have a relatively balanced score distribution in terms of technical quality, rationality, and naturalness.
Detailed score distribution of different categories in terms of gender. We also visualize the detailed score distribution of different categories for males and females in Fig. 15. We can observe that the score distribution of different categories for males and females is basically consistent except for the category of text-to-image. We speculate that people may have cognitive differences for familiar or unfamiliar objects, where women are more sensitive than men.
@@ 0.A.3 Detailed Score Distribution
Figure 17: Comparison of the original image and its corresponding low-frequency map.
Appendix 0.B Detailed Structure of the JOINT
@@ 0.B.1 Low-Frequency Map
0.B.2 Definition of Wasserstein Distance
@@ 0.B.3 Training Objective
Appendix 0.C More Experimental Details
0.C.1 Implementation Details
@@ 0.C.2 Evaluation Metrics
Figure 18: Success case of JOINT.
Figure 19: Failure case of JOINT.
0.C.3 Extended Qualitative Results
Success Cases. We visualized several successful cases (i.e., when the predicted results are consistent with subjective annotators in each dimension) for the proposed JOINT in Fig. 18. Specifically, we can observe that our JOINT is sensitive to global technical distortions such as blur, lack of detail, and contrast (1st and 2nd images in Fig. 18). Moreover, the perceptual artifacts-guided patch partition strategy endows JOINT with the ability to measure the severity of local artifacts (3rd and 4th images in Fig. 18). On the contrary, the rationality perceiving branch is not sensitive to technical distortion since we add deep feature regularization with filtered low-frequency information, and it is very sensitive to the rationality of contents and can recognize the very unusual compositions and objects in images (the left images in Fig. 18). All these cases further demonstrate the effectiveness of the JOINT in jointly learning from both perspectives in image naturalness assessment.
Failure Cases. As for the failure cases, we notice that most of them are difficult to recognize (poor sensory clarity, the right two in Fig. 19) or are extremely blurry with a small part of clear areas (mirror reflections, the left two in Fig. 19). Such cases are more frequent in photography that photographers gain prominence over the target by blurring the background, yet the model can experience challenges in predicting technical quality and rationality, resulting in a failure in naturalness evaluation.